OpenStudy (anonymous):

5y'' + 2y' +20y=cos(t) Find the steady state solution to the forced equation.

OpenStudy (anonymous):

Would the general solution to the equation be of any value?

OpenStudy (anonymous):

no i dont think so .. i think its just the steady state solution

OpenStudy (anonymous):

general is complementary + particular complementary is transient particular is steady state

OpenStudy (anonymous):

Plug in A Sin(t)+ B cos(t) into y and solve

OpenStudy (anonymous):

y=A sin(t)+ B cos(t) y'= A Cos(t)-B Sin(t) y''= -A sin(t)-B cos(t)

OpenStudy (anonymous):

5(-A sin(t)-B cos(t))+2(A cos(t)-B Sin(t))+20 (A sin(t)+B cos(t)) -5 A Sin(t)-5 cos(t)+ 2A cos(t)-2B Sin(t)+ 20 A sin(t)+20B cos(t)= cos(t)

OpenStudy (anonymous):

find A and b

OpenStudy (anonymous):

\[\frac{1}{229} (2 \sin (t)+15 \cos (t)) \]

OpenStudy (anonymous):

Of course, one has to do first what @imranmeah91 did above

OpenStudy (anonymous):

-5A-2B+20A=0 -5B+2A+20B=1